OpenStudy (anonymous):

Here is a ridiculously easy sample question from Oklahoma University's document on preparing for the Putnam Exam. Whoever solves first gets a medal. Took me about 3-4 minutes without a calculator =P. If you beat me on the time as well, then I'll fan you too :D "Find, with explanation, the maximum value of \(\bf f(x)=x^3-3x\) on the set of all real numbers \(\bf x\) satisfying \(\bf x^4+36 \le 13x^2\)."

OpenStudy (anonymous):

4:31 AM right now. Your time begins now. (You may also time yourself independently and please be honest lol)

OpenStudy (anonymous):

Also, please try to solve the question YOURSELF, i.e. without getting help from ANY other source.

OpenStudy (anonymous):

this is what you do for fun?

OpenStudy (anonymous):

lol

OpenStudy (dumbcow):

umm guess its 18 since domain is 2<x<3

OpenStudy (anonymous):

LOL @dumbcow got it. but he gets no medal since he guessed.

OpenStudy (dumbcow):

haha :)

OpenStudy (anonymous):

had he not included the word "guess" and i would've given him half a medal =]

OpenStudy (anonymous):

I should also include, legit solutions only.

OpenStudy (zzr0ck3r):

we know it has a max, but -1 and 1 does not satisfy the restraint, so solve the restraint and check the endpoints of the interval.

OpenStudy (dumbcow):

@zzr0ck3r will provide that for me

OpenStudy (zzr0ck3r):

im not good at solving those things

OpenStudy (zzr0ck3r):

lol

OpenStudy (anonymous):

@zzr0ck3r you're about to be a math major. stop kidding me.

OpenStudy (zzr0ck3r):

tell someone to let a = x^2 and that might help

OpenStudy (anonymous):

and for the odd powers, i don't think the squared substitution would help

OpenStudy (anonymous):

@zzr0ck3r did make a good suggestion which can be used in SOME part of the question which I will not state.

OpenStudy (anonymous):

basically re-arrange the inequality, so it comes down to finding the maximum of a function over some interval

OpenStudy (anonymous):

to lazy to work out details

OpenStudy (psymon):

Eh, you find the domain that was mentioned previously, in which you check your interval endpoints, get x = 3, plug it in and get 18 as the max value.

OpenStudy (dumbcow):

thats what i said

OpenStudy (campbell_st):

well looking at the inequality you get \[(x^2 -9)(x^2 -4)\le 0\]

OpenStudy (anonymous):

@Jack111 Everyone is telling me random stuff but noone has given a solution.

OpenStudy (psymon):

I know :P

OpenStudy (zzr0ck3r):

x^4 +36-13x^2≤0 (x^2)^2-13(x^2)+36<=0 (x^2-9)(x^2-4)<=0

OpenStudy (zzr0ck3r):

check 3,-3, 2, -2. one will be 18

OpenStudy (anonymous):

i am to lazy, if you know pretty much have the idea and the rest is just computation, then why bother

OpenStudy (zzr0ck3r):

if @dumbcow was right

OpenStudy (anonymous):

@zzr0ck3r stop helping unless you're ACTUALLY going to finish the solution lol.

OpenStudy (zzr0ck3r):

I think you can do this with Lagrange multipliers as well..

OpenStudy (anonymous):

The beauty of problem-solving is solving questions creatively with limited knowledge. That's how we gain and create more mathematical knowledge. In short, there is no need to use lagrange multipliers for a problem as simple as this.

OpenStudy (zzr0ck3r):

we know it has a max y'=0 gives x=-1,1 and these do not satisfy the restraint x^4 +36-13x^2≤0 (x^2)^2-13(x^2)+36<=0 (x^2-9)(x^2-4)<=0 this is cubic so it will be the "right end of the interval" f(3) = 18

OpenStudy (zzr0ck3r):

cubic with positive coeffecient

OpenStudy (zzr0ck3r):

ps of course I know there is no need:) and its pretty much the same thing anyway

OpenStudy (anonymous):

@zzr0ck3r You're solution is to an extent correct but still lacking.

OpenStudy (zzr0ck3r):

lol how so?

OpenStudy (zzr0ck3r):

and the domain would be [-3,-2] U [2,3]

OpenStudy (anonymous):

OK, I feel its complete enough now lol. @zzr0ck3r wins.

OpenStudy (zzr0ck3r):

I should have said its cubic with no square term

OpenStudy (anonymous):

$$x^4-13x^2+36\le 0\\(x^2-9)(x^2-4)\le 0$$hence for \([-3,-2]\cup[2,3]\) is the domain of optimization. for \(f(x)=x^3-3x\) observe:$$0=3(x^2-1)\\f''(x)=6x$$ i.e. \(x=\pm 1\) are our relative extrema. for \(x=-1\) we see \(f''<0\) i.e. negative curvature so we have a maximum at \(x=-1\) but this is not in our domain of optimization. since \(f\) is increasing an unbounded for \(x\gt1\) the answer is \(x=3\)

OpenStudy (anonymous):

not sure how long it took me probably 3-4 minutes

OpenStudy (zzr0ck3r):

you still didn't give the max value:P

OpenStudy (anonymous):

oh derp \(f(3)=3^3-3^2=18\)

OpenStudy (zzr0ck3r):

:)